OpenStudy (sydthekid913):

Which of the following points is a solution to the system of linear inequalities? 2x+3y>-15 -x-3y (greater than or equal to) -3

OpenStudy (sleepyjess):

What are the options?

OpenStudy (sydthekid913):

A. (3,-15) B. (3,-7) C. (3,0) D. (3,7)

OpenStudy (sleepyjess):

Just plug them in \(2(3)+3(-15)>-15\) \(-3-3(-15)\ge-3\) 2(3)+3(-7)>-15 \(-3-3(-7)\ge-3\) \(2(3)+3(0)>-15\) \(-3-3(0)\ge-3\) \(2(3)+3(7)>-15\) \(-3-3(7)\ge-3\)

OpenStudy (sydthekid913):

thank you sooooo much!!!

OpenStudy (sydthekid913):

\(\huge\ddot\smile\)

OpenStudy (sydthekid913):

how do you do the colored ones?

OpenStudy (sydthekid913):

I got it wrong

OpenStudy (sydthekid913):

I think I missed a sign or something:(